XYZ Company just paid a quarterly dividend of $3. The quarterly dividends grow at an effective
quarterly rate of 1.5%. XYZ’s effective quarterly cost of capital 5%. What is the expected price of
XYZ stock in 3 years?
Answer #1

Quarterly Rate = (1.015)1/4 - 1 = 0.37%

Quarterly Growth Rate = (1.05)1/4 - 1 = 1.23%

Dividend in 13th quarter = 3(1.0037)13= $3.15

Using DDM Model,

Stock Price in Year 3 = 3.15/(0.0123 - 0.0037)

Stock Price in Year 3 = $366.28